OpenStudy (anonymous):

If f(x) = int_{2}^{x} 1/{1+t^5} dt then f'(x) =

OpenStudy (danjs):

recall that the integral is also called an antiderivative

OpenStudy (danjs):

\[\frac{ d }{ dx }\int\limits x^2 dx = x^2 \]

OpenStudy (danjs):

oh is that an integral with bounds of 2 to x of 1/(1+t^5)dt

OpenStudy (anonymous):

yeah thats it

OpenStudy (jhannybean):

\[\int_2^x\frac{1}{1+t^5}dt\]

OpenStudy (perl):

d/dx integral {a,x} f(t) = f(x)

OpenStudy (jhannybean):

So the opposite of this is also true. \[d(\tan^{-1}(x)) = \frac{1}{1+x^2} \implies \int\frac{1}{1+x^2}dx = \tan^{-1}(x)\]

OpenStudy (jhannybean):

Sorry, I am missing something. \[\int\frac{du}{a^2+u^2} = \frac{1}{a}\tan^{-1}\left(\frac{u}{a}\right)+C\]

OpenStudy (perl):

you want to use the fundamental theorem of calculus here

OpenStudy (perl):